# Exact $n+2$ Comparison Complexity for the N-Repeated Element Problem

### v1

@@ 3 Adversary Argument
This form of argument is typically established through an adversary method. We imagine a two-player game between the algorithm and an adversary, similar to a chess match. At each step, the algorithm chooses two elements to compare, and the adversary must respond immediately with either “equal” or “not equal.” Eventually, the algorithm must commit to an answer—its guess for the repeated element. The adversary wins if it can produce a concrete input consistent with all previous answers yet for which the algorithm’s output is incorrect.
@@ 5 Finding a Large Clique
The key insight comes from viewing $P$ in terms of its complement $I$, the inequality graph. Let $C$ denote a minimum vertex cover of $I$. If the cover $C$ is small—that is, if |C|<n|C|<n—then the vertices outside the cover, $V(I)\setminus C$, induce a complete subgraph in $P$. Indeed, no edge of $I$ connects any two of these vertices, so all possible edges among them exist in $P$. Hence $P$ contains a clique of size
@@ 6 Finding Disjoint Cliques
Let $G$ have a minimum vertex cover of size exactly $k$. Then the edges of $G$ form a perfect matching: no two edges share a common vertex.
@@
Figure 1: Two disjoint $n$-cliques in $P$ from perfect matching in $I$. Blue pillars = “not equal” edges; green edges = cliques in $P$.

### v2

@@ Abstract
Keywords: comparison complexity; N-repeated element problem; equality comparison model; adversary argument; inequality graph; lower bound.
@@ 3 Adversary Argument
This form of argument is typically established through an adversary method [2]. We imagine a two-player game between the algorithm and an adversary, similar to a chess match. At each step, the algorithm chooses two elements to compare, and the adversary must respond immediately with either “equal” or “not equal.” Eventually, the algorithm must commit to an answer—its guess for the repeated element. The adversary wins if it can produce a concrete input consistent with all previous answers yet for which the algorithm’s output is incorrect.
@@ 5 Finding a Large Clique
The key insight comes from viewing $P$ in terms of its complement $I$, the inequality graph. Let $C$ denote a minimum vertex cover [3] of $I$. If the cover $C$ is small—that is, if |C|<n|C|<n—then the vertices outside the cover, $V(I)\setminus C$, induce a complete subgraph in $P$. Indeed, no edge of $I$ connects any two of these vertices, so all possible edges among them exist in $P$. Hence $P$ contains a clique of size
@@ 6 Finding Disjoint Cliques
Let $G$ have a minimum vertex cover of size exactly $k$. Then the edges of $G$ form a perfect matching [4]: no two edges share a common vertex.
@@
Figure 1: Two disjoint $n$-cliques in $P$ from a perfect matching in $I$. Blue pillars represent “not equal” edges; green edges represent cliques in $P$.
@@ 8 Discussion
Acknowledgements
This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.